# Danny Halawi4, Fred Zhang4, Philip E. Tetlock2,5

### v4

@@ Abstract
footnotetext: Preprint.footnotetext: * Equal Contributionfootnotetext: Correspondence to benchmark@forecastbench.org.footnotetext: This work was funded by a grant from Open Philanthropy; the grant supports the maintenance and updating of this benchmark for three years. We thank Sam Glover and Molly Hickman for writing superforecaster prompts, Otto Kuusela for helpful discussions, and Harrison Durland and Victoria Schmidt for their research assistance. The views expressed in this paper do not necessarily reflect the views of the Federal Reserve Bank of Chicago or the Federal Reserve system.
@@ 1 Introduction
Forecasting the future is a challenging but important task (Armstrong, 2001, Tetlock and Gardner, 2015). Economic forecasts influence investment and hiring decisions (Christensen et al., 2018). And forecasts of the Covid-19 pandemic in early 2020 prompted local lockdowns to slow the spread of the virus Adam (2020). However, human forecasting is often expensive, time-consuming, applicable only in specific domains, and subject to concerns about human biases. Motivated by these limitations, recent work explores the use of machine learning (ML) models, especially large language models (LLMs), in automated forecasting (Hendrycks et al., 2021, Zou et al., 2022, Yan et al., 2024, Pratt et al., 2024, Phan et al., 2024, Fluri et al., 2024, Halawi et al., 2024).
To assess LLMs’ forecasting capabilities, prior work built static benchmarks of questions where the answer was known (resolved) after the knowledge cut-offs of the LLMs they test (Zou et al., 2022, Yan et al., 2024, Halawi et al., 2024). For example, “Will a nuclear weapon be detonated in 2023,” though resolved on Jan 1, 2024, is a valid out-of-sample question for testing a model with a known knowledge cut-off before the end of 2023.
This static evaluation methodology comes with three key drawbacks. First, as the knowledge cut-offs of frontier models are updated, static benchmarks become obsolete, necessitating further data-sourcing. This makes it difficult to continuously track and compare the top models in the field. Second, knowledge cut-offs are usually estimated using the time range of pre-training data. Such estimates may not be accurate, and post-training may inject further post-cutoff knowledge into the model, contaminating the test sets. Lastly, model developers face financial incentives to exaggerate their accuracy on benchmarks and claim that their models are state-of-the-art performers. While some fudging is easily identified, other subtle benchmark manipulation or overfitting is harder to catch, and some studies have shown that there is significant evidence for benchmark contamination and/or memorization in popular LLM models (Roberts et al., 2023, Elazar et al., 2024, Li et al., 2023).
@@
Our initial benchmark consists of 1,000 standardized forecasting questions randomly sampled from a much larger real-time question bank. To establish baseline levels of performance, we record predictions from dozens of LLMs on the initial set, using methods like retrieval-augmentation to boost model performance (Lewis et al., 2020, Halawi et al., 2024). We independently elicit predictions from human forecasters of varying expertise, including a group of seasoned forecasters (superforecasters) (Tetlock and Gardner, 2015) who have consistently performed well in competitive forecasting tournaments. As questions resolve, we rate the submissions of registered models and the human comparison groups, updating our public leaderboard.
Our initial results indicate that state-of-the-art models, such as GPT-4o and Claude-3.5, perform only roughly as well as a simple median of forecasts from a survey of humans with no (or minimal) forecasting experience, even when the LLMs are augmented with news retrieval and prompt engineering, and when the models have access to forecasts from a human crowd (on market-based questions). The models also perform significantly worse than the median of expert forecasters (superforecasters) who answer the same questions.
These findings leave significant room for researchers to attempt to improve AI-based forecasting systems using innovative approaches, such as developing methods for continuously updating models with current events and enhancing LLMs to reason over extended time frames. To support progress in this area, we publish an auxiliary dataset of LLM and human forecasts, rationales, and accuracy for use in future LLM fine-tuning and testing.
@@ 1.1 Related work
ML systems that make accurate forecasts can help inform human decision-making (Hendrycks et al., 2021, Schoenegger et al., 2024a). Recent work studies the use of LLMs for automated forecasting Jin et al. (2021), Zou et al. (2022), Yan et al. (2024), Halawi et al. (2024), Pratt et al. (2024). These papers all build static benchmarks of resolved questions. A recent paper from Halawi et al. (2024) uses questions resolved between June 2023 and January 2024. Unfortunately, the latest LLMs, such as GPT-4o, have knowledge cut-offs past this point. This fact motivates our work to build a dynamically updating benchmark that can accurately evaluate frontier model performance.
In addition, Schoenegger and Park (2023) and Abolghasemi et al. (2023) compare the accuracy of GPT-4 and other LLMs to human forecasters. Schoenegger et al. (2024b) found that an ensemble of $12$ LLMs rivaled human forecasts in a forecasting tournament in 2023 with a small number of questions, limiting the study’s statistical power. Unlike our work, that tournament was run only once and is no longer updated. Also, our much larger question set allows us to make precise statistical claims about the performance of LLMs relative to each other and relative to human performers.
Finally, recent work focuses on the use of LLMs and transformer-based systems in statistical time-series forecasting (Nie et al., 2023, Gruver et al., 2023, Dooley et al., 2023, Rasul et al., 2023, Jin et al., 2024, Das et al., 2024, Goswami et al., 2024, Woo et al., 2024), but many of the most important forecasting questions do not have well-defined time series that can be used in standard statistical forecasting models (e.g., what is the probability that a nuclear weapon will be used offensively in the next ten years?). Our benchmark is more general, and evaluates automated forecasting performance across questions with and without underlying time series and historical baseline data.
@@
Evaluating highly capable LLMs is a challenging task—with many models saturating key benchmarks soon after their release (Maslej et al., 2023, Owen, 2024) and with benchmarks potentially leaked to models’ training data (Magar and Schwartz, 2022, Sainz et al., 2023, Jacovi et al., 2023, Xu et al., 2024b, Balloccu et al., 2024, Jiang et al., 2024b, Zhang et al., 2024, Xu et al., 2024a). Our dynamic forecasting benchmark avoids both of these issues. First, automated forecasting is highly unsaturated: Halawi et al. (2024) showed that under simple zero-shot evaluation, frontier models such as GPT-4 are much less accurate than aggregates of human predictions. Second, our benchmark is dynamic. The final resolution of each question is only determined in the future and cannot be ‘leaked’ in any training data (by construction). This eliminates concerns of contamination.
@@ 2 Preliminaries
We evaluate 17 LLMs on our initial benchmark: GPT-3.5-Turbo-Instruct (Brown et al., 2020), GPT-4 (OpenAI, 2023), GPT-4o, Llama-2-70B (Touvron et al., 2023), Llama-3-7B, Llama-3-70B, Mistral-7B (Jiang et al., 2023), Mixtral-8x7B (Jiang et al., 2024a), Mixtral-8x22B, Mistral-Large, Qwen1.5-110B-Chat (Bai et al., 2023), Claude-2.1 (Anthropic, 2023), Claude-3-Haiku, Claude-3.5-Sonnet, Claude-3-Opus (Anthropic, 2024), Gemini 1.5 Flash and Gemini 1.5 Pro (Gemini Team, 2023).
We record model predictions under three different methods: zero-shot prompting, scratchpad prompting, and scratchpad prompting with retrieval augmentation. For the last setting, we use the retrieval infrastructure from Halawi et al. (2024) and provide relevant news articles to the models in-context to reason about. Additionally, only models with a context window larger than 8,000 tokens were evaluated under the retrieval setting due to the inclusion of news articles in some prompts.
@@ 3 Benchmark, leaderboard, and datasets
Every night, our automated system updates the question bank with new questions and resolution values. We drop invalid, low-quality, and resolved questions, categorizing the remaining questions by topic. The process is fully automated as detailed in Section 3.1.
Every two weeks, we sample from the question bank and release question sets for teams interested in submitting their forecasts to the benchmark. We also survey a standard set of LLM-based models to allow for comparisons of performance over time. The forecasts that these groups submit are resolved continuously with daily updates to our leaderboard. We provide the resulting data on forecast questions and submissions to researchers. See Section 3.2 for details. Finally, we discuss the question resolution procedure in Section 3.3 and our leaderboard design in Section 3.4.
@@ 3.1.1 Questions and resolution values
As one of our two main question sources, we rely on a handful of prediction markets and forecast aggregation sites that elicit human predictions on questions across a wide range of topics. [^2] Hereafter, for brevity, we refer to questions that come from both prediction markets and forecast aggregation sites as ”market” questions. We also refer to theses sources collectively as ”market” sources. When selecting questions from market sources to add to the question bank, we choose those with high levels of active human trading on these platforms (liquidity) as these questions tend to be of higher quality than those with low levels of human trading. [^3] See Table 8 for an example question pulled from a market source. We also store the latest market and resolution values for each question.
@@
As a second question source, we rely on 5 well-established and well-maintained datasets that track real-world events (e.g. ACLED (Raleigh et al., 2023), a geopolitical database that tracks worldwide conflict, including facts like the number of protests in Niger each month). With these dataset sources, we can generate questions using a fixed question template (e.g., “Will the number of protests in Niger increase by at least 10% over this month’s value by the resolution date?"). [^4] See Table 9 for an example question pulled from a dataset source.
@@
RFI randforecastinginitiative.org 15 105
Manifold Markets manifold.markets 392 76,636
Metaculus metaculus.com 800 319,600
Polymarket polymarket.com 534 142,311
Market Total 1,741 538,652
ACLED acleddata.com 3,150 4,959,675
@@
Wikipedia wikipedia.org 335 55,945
Yahoo! Finance finance.yahoo.com 504 126,756
Dataset Total 4,207 5,157,397
Question Bank Total 5,948 5,696,049
@@ 3.1.2 Question metadata
After we automatically collect forecasting questions and resolution values in our standardized question bank, the data get processed in several ways. This generates more information about the questions and creates another sampling option for the creation of the final question set. Following Halawi et al. (2024), we use gpt-3.5-turbo-0125 to categorize questions by subject and to filter out low-quality questions. We report the number of questions by category and source in Table 15, where we display the breakdown of the standard questions from our question bank ($N$ from Table 1) across question categories by source, after removing low-quality questions.
@@ 3.2 Question sets
We release a set of 1,000 forecast questions for LLMs every other Sunday at midnight UTC. We sample an equal number of questions from each data source to ensure representativeness. Within each source, we then uniformly sample questions across all question categories, aiming for an equal distribution from each category within each source. This ensures that models cannot be overfit to a specific type of question or topic. Limiting the number of questions generated to 1,000 also ensures that costs for testing LLM systems on the benchmark are capped for development teams with fewer resources.
@@
The human question is comprised of 200 forecast questions sampled directly from the LLM question set. When sampling from the LLM question set, we do our best to maintain proportionality across question sources and across categories within each question source; this ensures the question set addresses as many domains as possible.
@@
For questions derived from dataset sources, the distribution of resolution dates is the $forecast\_due\_date+n$ days, where $n\in\{7,30,90,180,365,1095,1825,3650\}$. In other words, for each dataset question we ask for 8 forecasts, differing only in their resolution date, ranging from one week to 10 years. For questions derived from market sources, we ask for only 1 forecast: the probability that each question will resolve positively (will the event underlying the question occur, or not). In time, this setup will allow us to evaluate both human and LLM forecasting performance over the short, medium, and long term.
@@
For every data source, there are two types of questions, each comprising half of the question set. The first type is a standard forecasting question with a binary outcome, e.g. “Will inflation (core CPI) be above 3% next month?” We construct the second type, combination questions, by pairing two standard questions. For combination questions, we ask for forecasts on all Boolean combinations of the two questions (i.e. $P(Q1\cap Q2)$, $P(Q1\cap\neg Q2)$, …). Considering the extensive number of existing standard questions and potential combinations in our question bank (as we could potentially combine 3, 4, or more standard questions together), we effectively have access to millions of possible forecasting questions to sample from. We show the number of two-question combination questions in the question bank as it stands at time of writing in the right-hand column of Table 1. This setup implies that for market combination questions, each forecaster will provide 4 forecasts, whereas for dataset combination questions, each forecaster will provide 32 forecasts (4 for each boolean combination of $Q1$ and $Q2$ at each of the 8 forecast horizons).
@@
Of the 1,000 questions in the LLM question set, 500 are combination questions. Each combination question is composed of two standard questions from the same set. This means LLMs will also provide forecasts for the individual components of these combination questions, since they’re drawn from the existing standard questions. Importantly, none of the 200 questions in the human question set are combination questions. [^5] We exclude combination questions because expert human forecasts are expensive to generate and by focusing on forecasts of standard questions we maximize the relevance of these expert human forecasts.
@@
To estimate baseline human and expert performance on these questions, we run surveys asking the general public and expert forecasters to predict on our question sets (see Section 4). We produce the question sets 10 days before the forecast due date to allow for time to create and run a human survey. LLM teams receive their question set 24 hours before the due date, even though it was generated at the same time as the human question set; this constrained time frame gives teams less time to be able to game the system, or be dishonest participants on the leaderbord. We thus elicit forecasts, obtaining comparable forecast sets on the due date from both LLMs and humans who faced the same information environment.
@@ 3.3 Resolution
We resolve forecast sets nightly by gathering the latest information about which events have or have not occurred. All questions are ultimately resolved to the final true value.
@@
Ground truth is not available for market questions before the events are known to have occurred or not occurred. Before questions resolve, we evaluate performance by calculating the squared distance between the forecast value and the platform prediction (an aggregate of human forecasts reported on the platform) from the preceding day. These values are updated nightly as described in Section 3.1.1. This allows us to score the complete set of forecast questions, incorporating all information available to people on a nightly basis. Once a market question is officially resolved, we score the forecast against the resolution value, creating a definitive score for the question.
Evaluating forecasts on unresolved market questions to the crowd forecast provides a good estimate of forecast performance, as crowd forecasts tend toward ground truth as the resolution date approaches. This allows us to estimate forecast performance on a larger set of market questions, as we would otherwise have to wait until each market resolved.
@@
Datasets can be updated as new information becomes available. Thus, questions derived from datasets are continuously resolved to the value from the the latest available data. As the resolution dates for dataset questions mentioned in earlier come due (the 7 day horizon to the 10 year horizon), a new round of forecasts is evaluated. We thus are able to evaluate forecasting performance over different time horizons.
@@
We select 1,000 questions for the LLM question set to make the forecasting task impractical to complete manually within the 24 hour window after the question set is released. And we obligate all LLMs to forecast on all questions to ensure comparability of scores across models and human-based aggregates. So, when a model does not submit predictions on certain questions or time horizons, we consider that a model error and impute a naive value for the model to ensure comparability across models over time.
For market questions, as we only ask users to forecast the outcome of the question, missing forecasts on question outcomes are assigned the value of the crowd forecast on the forecast due date. Some may argue this is overly-generous, but we did not want teams to have a competitive advantage by simply scraping the market websites themselves.
For dataset questions, we impute the value $0.5$ (which represents a lack of information) to forecasts across all time horizons. Empirically, top models report valid forecasts on all questions and are not affected by this imputation procedure.

### v5

# Ezra Karger

@@ Abstract
[…]
@@ 1 Introduction
Forecasting the future is a challenging but important task (Armstrong, 2001; Tetlock and Gardner, 2015). Economic forecasts influence investment and hiring decisions (Christensen et al., 2018). And forecasts of the Covid-19 pandemic in early 2020 prompted local lockdowns to slow the spread of the virus (Adam, 2020). However, human forecasting is often expensive, time-consuming, applicable only in specific domains, and subject to concerns about human biases. Motivated by these limitations, recent work explores the use of machine learning (ML) models, especially large language models (LLMs), in automated forecasting (Hendrycks et al., 2021; Zou et al., 2022; Yan et al., 2024; Pratt et al., 2024; Phan et al., 2024; Fluri et al., 2024; Halawi et al., 2024).
To assess LLMs’ forecasting capabilities, prior work built static benchmarks of questions where the answer was known (resolved) after the knowledge cut-offs of the LLMs they test (Zou et al., 2022; Yan et al., 2024; Halawi et al., 2024). For example, “Will a nuclear weapon be detonated in 2023,” though resolved on Jan 1, 2024, is a valid out-of-sample question for testing a model with a known knowledge cut-off before the end of 2023.
This static evaluation methodology comes with three key drawbacks. First, as the knowledge cut-offs of frontier models are updated, static benchmarks become obsolete, necessitating further data-sourcing. This makes it difficult to continuously track and compare the top models in the field. Second, knowledge cut-offs are usually estimated using the time range of pre-training data. Such estimates may not be accurate, and post-training may inject further post-cutoff knowledge into the model, contaminating the test sets. Lastly, model developers face financial incentives to exaggerate their accuracy on benchmarks and claim that their models are state-of-the-art performers. While some fudging is easily identified, other subtle benchmark manipulation or overfitting is harder to catch, and some studies show significant evidence of benchmark contamination and/or memorization by popular LLM models (Roberts et al., 2023; Elazar et al., 2024; Li et al., 2023).
@@
Our initial benchmark consists of 1,000 standardized forecasting questions randomly sampled from a much larger real-time question bank. To establish baseline levels of performance, we record predictions from dozens of LLMs on the initial set, using methods like retrieval-augmentation to boost performance (Lewis et al., 2020; Halawi et al., 2024). We independently elicit predictions from two groups of human forecasters: the general public and seasoned forecasters (superforecasters) (Tetlock and Gardner, 2015) who have consistently performed well in competitive forecasting tournaments. As questions resolve, we rate the submissions of registered models and the human comparison groups, updating our public leaderboard.
Our initial results indicate that state-of-the-art models, such as Claude-3.5 Sonnet and GPT-4 Turbo, perform only roughly as well as a simple median of forecasts from a survey of humans with no (or minimal) forecasting experience, even when the LLMs are augmented with news retrieval and prompt engineering and when they have access to forecasts from a human crowd (on market-based questions). The models also perform significantly worse than the median forecast of superforecasters.
These findings leave significant room for researchers to improve AI-based forecasting systems using innovative approaches, such as developing methods for continuously updating models with current events and enhancing LLMs to reason over extended time frames. To support progress in this area, we publish an auxiliary dataset of LLM and human forecasts, rationales, and accuracy for use in future LLM fine-tuning and testing.
@@ 1.1 Related work
ML systems that make accurate forecasts can help inform human decision-making (Hendrycks et al., 2021; Schoenegger et al., 2024a). Recent work studies the use of LLMs for automated forecasting (Jin et al., 2021; Zou et al., 2022; Yan et al., 2024; Halawi et al., 2024; Pratt et al., 2024). These papers all build static benchmarks of resolved questions. A recent paper from Halawi et al. (2024) uses questions resolved between June 2023 and January 2024. Unfortunately, the latest LLMs have knowledge cut-offs past this point. This fact motivates our work to build a dynamically updating benchmark that can accurately evaluate frontier model performance.
In addition, Schoenegger and Park (2023) and Abolghasemi et al. (2023) compare the accuracy of GPT-4 and other LLMs to human forecasters. Schoenegger et al. (2024b) found that an ensemble of $12$ LLMs rivaled human forecasts in a forecasting tournament with a small number of questions, limiting the study’s statistical power. Unlike our work, that tournament was run only once and is no longer updated. Also, our much larger question set allows us to make precise statistical claims about the performance of LLMs relative to each other and to humans.
Finally, recent work focuses on the use of LLMs and transformer-based systems in statistical time-series forecasting (Nie et al., 2023; Gruver et al., 2023; Dooley et al., 2023; Rasul et al., 2023; Jin et al., 2024; Das et al., 2024; Goswami et al., 2024; Woo et al., 2024), but many of the most important forecasting questions do not have well-defined time series that can be used in standard statistical forecasting models (e.g., what is the probability that a nuclear weapon will be used offensively in the next ten years?). Our benchmark is more general, and evaluates forecasting performance across questions with and without underlying time series and historical baseline data.
@@
Evaluating highly capable LLMs is a challenging task—with many models saturating key benchmarks soon after their release (Maslej et al., 2023; Owen, 2024) and with benchmarks potentially leaked to models’ training data (Magar and Schwartz, 2022; Sainz et al., 2023; Jacovi et al., 2023; Xu et al., 2024b; Balloccu et al., 2024; Jiang et al., 2024b; Zhang et al., 2024; Xu et al., 2024a). Our dynamic forecasting benchmark avoids both of these issues. First, automated forecasting is highly unsaturated: Halawi et al. (2024) showed that under simple zero-shot evaluation, frontier models such as GPT-4 are much less accurate than aggregates of human predictions. Second, our benchmark is dynamic. The final resolution of each question is only determined in the future and cannot be leaked in any training data (by construction). This eliminates concerns of contamination.
@@ 2 Preliminaries
We evaluate 17 LLMs on our initial benchmark: GPT-3.5-Turbo-Instruct (Brown et al., 2020), GPT-4 (OpenAI, 2023), GPT-4o, Llama-2-70B (Touvron et al., 2023), Llama-3-7B, Llama-3-70B, Mistral-7B (Jiang et al., 2023), Mixtral-8x7B (Jiang et al., 2024a), Mixtral-8x22B, Mistral-Large, Qwen1.5-110B-Chat (Bai et al., 2023), Claude-2.1 (Anthropic, 2023), Claude-3-Haiku, Claude-3.5-Sonnet, Claude-3-Opus (Anthropic, 2024), Gemini 1.5 Flash, and Gemini 1.5 Pro (Gemini Team, 2023). We outline the various baselines we run with these models in Section 5.
@@ 3 Benchmark, leaderboard, and datasets
Every night, our automated system updates the question bank with new questions and resolution values. We drop invalid, low-quality, and resolved questions, categorizing the remaining questions by topic. The process is fully automated, as detailed in Section 3.1.
Every two weeks, we sample from the question bank and release question sets for those interested in submitting their forecasts to the benchmark. We also survey a standard set of LLM-based models to allow for comparisons of performance over time. Submitted forecasts are resolved continuously with daily updates to our leaderboard. We provide the resulting data on forecast questions and submissions to researchers. See Section 3.2 for details. Finally, we discuss the question resolution procedure in Section 3.3 and our leaderboard design in Section 3.4.
@@ 3.1.1 Questions and resolution values
We compile market questions from a handful of prediction markets and forecast aggregation sites that elicit human predictions on questions across a wide range of topics. [^2] Hereafter, for brevity, we refer to questions that come from both prediction markets and forecast aggregation sites as “market” questions. We also refer to theses sources collectively as “market” sources. When selecting questions from market sources to add to the question bank, we choose those with high levels of active human trading on these platforms (liquidity) as these questions tend to be of higher quality than those with low levels of human trading. [^3] See Table 8 for an example question pulled from a market source. We store the latest market and resolution values for each question.
@@
We create dataset questions from well-established and well-maintained datasets that track real-world events (e.g., ACLED (Raleigh et al., 2023), a geopolitical database that tracks worldwide conflict, including facts like the number of protests in Niger each month). With these sources, we can generate questions using a fixed question template (e.g., “Will the number of protests in Niger increase by at least 10% over this month’s value by the resolution date?”). [^4] See Table 9 for an example question pulled from a dataset source.
@@
RFI randforecastinginitiative.org 18 153
Manifold Markets manifold.markets 405 81,810
Metaculus metaculus.com 722 260,281
Polymarket polymarket.com 915 418,155
Market Total 2,060 760,399
ACLED acleddata.com 3,220 5,182,590
@@
Wikipedia wikipedia.org 428 91,378
Yahoo! Finance finance.yahoo.com 509 129,286
Dataset Total 4,375 5,418,275
Question Bank Total 6,435 6,178,674
@@ 3.1.2 Question metadata
After we automatically update the question bank with new forecasting questions and resolution values, the data are processed in several ways. This generates more information about the questions and creates another sampling option when creating the question set. Following Halawi et al. (2024), we use gpt-4o-mini to categorize questions by subject and to filter out low-quality questions. We report the number of questions by category and source in Table 15, where we display the breakdown of the standard questions from our question bank ($N$ from Table 1) across question categories by source, after removing low-quality questions.
@@ 3.2 Question sets
We release a set of 1,000 forecast questions for LLMs every other Sunday at midnight UTC. We sample an equal number of questions from each source to ensure representativeness. Within each source, we then uniformly sample questions across all question categories, aiming for an equal distribution from each category within each source. This ensures that models cannot be overfit to a specific type of question or topic. Limiting the number of questions generated to 1,000 also ensures that costs for testing LLMs on the benchmark are capped for development teams with fewer resources.
@@
The human question set is comprised of 200 forecast questions sampled directly from the LLM question set. When sampling, we do our best to maintain proportionality across question sources and across categories within each question source; this ensures the question set addresses as many domains as possible.
@@
For questions derived from dataset sources, the distribution of resolution dates is the $forecast\_due\_date+n$ days, where $n\in\{7,30,90,180,365,1095,1825,3650\}$. In other words, for each dataset question we ask for 8 forecasts that differ only in their resolution date. For questions derived from market sources, we ask for only 1 forecast: the probability that each question will resolve positively (will the event underlying the question occur, or not). This setup will allow us to evaluate both human and LLM forecasting performance over the short, medium, and long term.
@@
There are two types of questions, each comprising half of the question set. The first type is a standard forecasting question with a binary outcome, e.g., “Will inflation (core CPI) be above 3% next month?” We construct the second type, combination questions, by pairing two standard questions. For combination questions, we ask for forecasts on all Boolean combinations of the two questions (i.e., $P(Q1\cap Q2)$, $P(Q1\cap\neg Q2)$, …). Considering the extensive number of standard questions and potential combinations in our question bank (as we could potentially combine 3, 4, or more standard questions together), we effectively have access to millions of possible forecasting questions from which we can sample. We show the number of two-question combination questions in the question bank as it stands at time of writing in the right-hand column of Table 1. This setup implies that for market combination questions, each forecaster will provide 4 forecasts, whereas for dataset combination questions, each forecaster will provide 32 forecasts (4 for each Boolean combination of $Q1$ and $Q2$ at each of the 8 forecast horizons).
@@
Of the 1,000 questions in the LLM question set, 500 are combination questions. Each combination question is composed of two standard questions from the same question set. This means that LLMs will also provide forecasts for the individual components of these combination questions, since they’re drawn from the existing standard questions. Importantly, none of the 200 questions in the human question set are combination questions. [^5] We exclude combination questions because expert human forecasters are expensive; we maximize their relevance by having them focus on standard questions.
@@
To compare human performance to LLMs, we periodically run surveys asking the general public and superforecasters to forecast on our question sets (see Section 4). We produce the question sets 10 days before the forecast due date to allow for time to create and run a human survey. LLM teams receive their question set 24 hours before the due date, even though it was generated at the same time as the human question set. This constrained time frame gives teams less time to be able to game the system. We thus elicit forecasts, obtaining comparable forecast sets on the due date from both LLMs and humans who faced the same information environment.
@@ 3.3 Resolution
We resolve forecast sets nightly by gathering the latest information about which events have or have not occurred. All questions are ultimately resolved to ground truth.
@@
For market questions, ground truth is not available until the question has been resolved on the platform. Until then, we evaluate performance by calculating the squared distance between the forecasted value and the platform’s crowd forecast (an aggregate of human forecasts reported on the platform) from the preceding day. [^6] The crowd forecasts are updated nightly in the Question Bank as described in Section 3.1.1. This provides a good estimate of forecast performance on unresolved questions since crowd forecasts tend toward ground truth as the resolution date approaches.
Once a market question has officially been resolved, we score the forecast against the resolution value, creating a definitive score for the question. We are thus able to estimate forecast performance on the entire set of market questions (resolved and unresolved), incorporating all information available to markets on a nightly basis.
@@
Datasets can be updated as new information becomes available. Hence, questions derived from datasets are continuously resolved to the value from the latest available data. As the resolution dates (ranging from a 7-day to a 10-year horizon) for dataset questions come due, a new round of forecasts is evaluated. We thus are able to evaluate forecasting performance over different time horizons.
@@
We select 1,000 questions for the LLM question set to make the forecasting task impractical to complete manually within the 24-hour window after the question set is released. And we obligate all LLMs to forecast on all questions to ensure comparability of scores across models and human-based aggregates. When a model does not submit forecasts on certain questions or time horizons, we consider that a model error and impute a naive forecast for the model to ensure comparability across models over time.
For market questions, since we only ask for forecasts on the outcome of the question, missing forecasts are assigned the value of the crowd forecast on the forecast due date. Some may argue this is overly-generous, but we did not want forecasters to have a competitive advantage by simply scraping the market websites themselves.
For dataset questions, we impute the value $0.5$ (which represents an uninformed 50% forecast) to forecasts across all time horizons. Empirically, top models report valid forecasts on all questions and are not affected by this imputation procedure.